1. Find the general solution by substitution or elimination. (p) -2 + y + 2e+ x - y-et
Answer #1

» Ye = ez+G02 Let 7p = Bet * DYpa ob et Dyp = Bet da Darn-y = rêt - cd -ét ano dy - et Dy axty- It - key = y het ш) DX 0 » Dhci-cē 2 got Yo = -zet siyat) YmYp (vinj >/y (4 egentzett in W,+ lising (wi) & (vi) f (61+₂ e 2t + ēt) - (e3+40 2t zet)=zet (~## Dont FORGET to give RATINGS.